True or false with full explanation


i. Every continuous function is differentiable.


ii. Every integrable function is monotonic

1
Expert's answer
2022-06-01T03:36:05-0400

ANSWER Both statements are false.

EXPLANATION

Let

"f(x)=\begin{cases}
 & \text -2 x+1,\, \, \, { if } \, -1\leq x<0 \\ 
 & \text0.4 x+1,\, \, { if } \, \,\, \, 0\leq x\leq 1
\end{cases}"

Since the function coincides with the polynomials on the segments "[-1,0)" and "(0,1]" then the function is continuous on these segments.

"\lim_{x\rightarrow 0^{-}}f(x)= \lim_{x\rightarrow 0^{-}}(-2x+1)=1=f(0) =\\=\lim_{x\rightarrow 0^{+}}f(x)= \lim_{x\rightarrow 0^{+}}(0.4x+1 )=1 ."

Thus, "f" is continuous at the point "x=0" and on the segment "[-1,1]."



(i)

The left-hand derivative at "x=0" is

"f _{-}^{'}(0)=\lim_{x\rightarrow 0^{-}}\frac{f(x)-f(1) }{x}=\lim_{x\rightarrow 0^{-}}\frac{-2x+1-1 }{x}=-2" ,

the right-hand derivative at "x=0" is

"f _{+}^{'}(0)=\lim_{x\rightarrow 0^{+}}\frac{f(x)-f(1) }{x}=\lim_{x\rightarrow 0^{+}}\frac{0.4x+1-1 }{x}=0.4"

Since "f_{-}^{'}(0)\neq f_{+}^{'}(0)" , then function is not differentiable at the point "x=0"

(ii) Because function "f" is continuous on the segment "[-1,1]" , then the function is integrable

on "[-1,1]." But this function is not monotonic on the segment "[-1,1]" , since in the interval "(-1,0)" the function decreases , and in the interval "(0,1)" the function increases .